OpenStudy (anonymous):

f(x) = \frac{1}{x - 3} Use the limit definition of the derivative on page 156 to find (i) f'( -1 ) = (ii) f'( 1 ) = (iii) f'( 5 ) = (iv) f'( 7 ) = To avoid calculating four separate limits, I suggest that you evaluate the derivative at the point when x = a. Once you have the derivative, you can just plug in those four values for "a" to get the answers.

OpenStudy (anonymous):

\[f(a)=\frac{1}{a-3}~~\Rightarrow~~f'(a)=\lim_{h\to0}\frac{f(a+h)-f(a)}{h}\] \[\begin{align*}f'(a)&=\lim_{h\to0}\frac{\dfrac{1}{a+h-3}-\dfrac{1}{a-3}}{h}\\\\ &=\lim_{h\to0}\frac{\dfrac{a-3}{(a+h-3)(a-3)}-\dfrac{a+h-3}{(a-3)(a+h-3)}}{h}\\\\ &=\lim_{h\to0}\frac{\dfrac{-h}{(a+h-3)(a-3)}}{h}\\\\ &=\lim_{h\to0}\frac{-1}{(a+h-3)(a-3)}\\\\\\ &=\cdots \end{align*}\]

OpenStudy (anonymous):

So you'll get some function of \(a\). To find (a)-(d), all you have to do is plug in the given \(a\).

OpenStudy (anonymous):

Sorry, I mean (i)-(iv)...

OpenStudy (anonymous):

thank you so much! will my answers include "h" in them or no?

OpenStudy (anonymous):

never mind i got it now thank you so much again!!

OpenStudy (anonymous):

You're welcome!